### v1

@@ 3.3 Variance penalty for MMD loss
$=\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}^{2}\left\langle\frac{\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}}{\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}},f\right\rangle_{\mathcal{H}}^{2}\right],$
where $\phi_{\boldsymbol{x}}=k(\boldsymbol{x},\cdot)$ and $\phi_{\boldsymbol{y}}=k(\boldsymbol{y},\cdot)$. With Cauchy inequality, it gives
$\left\langle\frac{\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}}{\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}},f\right\rangle_{\mathcal{H}}\leq\frac{\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}}{\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}}\|f\|_{\mathcal{H}}\leq 1.$
When $f=\frac{\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}}{\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}}\in\mathcal{F}$, the equality holds. Thus
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\boldsymbol{x})-f(\boldsymbol{y})\right]^{2}=\mathbb{E}\left[\|\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}}\|_{\mathcal{H}}^{2}\right].$
Then the second term of Eq. 21 can be rewritten as
(22) $\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\boldsymbol{x})-f(\boldsymbol{y})\right]^{2}=\mathbb{E}_{\boldsymbol{x}\sim p}[k(\boldsymbol{x},\boldsymbol{x})]-2\mathbb{E}_{\boldsymbol{x}\sim p,\boldsymbol{y}\sim q}[k(\boldsymbol{x},\boldsymbol{y})]+\mathbb{E}_{\boldsymbol{y}\sim p}[k(\boldsymbol{y},\boldsymbol{y})].$
Similar to Eq. 19, we can estimate Eq. 22 with ensemble particles, i.e.,
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\boldsymbol{x})-f(\boldsymbol{y})\right]^{2}=\sum_{i=1}^{m}$ $w_{i}k(\hat{\boldsymbol{x}}_{k}^{(i)},\hat{\boldsymbol{x}}_{k}^{(i)})-2\sum_{i,j=1}^{m}w_{i}v_{j}k(\hat{\boldsymbol{x}}_{k}^{(i)},\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(j)}))$
$+\sum_{i}^{n}v_{i}k(\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(i)}),\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(i)})),$
where $\{v_{i}=1/N\}_{i=1}^{N}$ are weights of particles generated by transport map $\mathcal{T}_{k}$ and $\{w_{i}\}_{i=1}^{N}$ are weights of particle filter defined in Eq. 18.
@@
If we restrict the kernel function to be the linear kernel and the search space of the transport map to be linear, we can obtain an analytical form of the transport map filter.
@@
Given the RKHS kernel $k(\boldsymbol{x},\boldsymbol{x}^{\prime})=\boldsymbol{x}^{\top}\boldsymbol{x}^{\prime}+1$ and regularization parameter $\lambda=1$. Moreover, we assume Eq. 16 to be a linear transport, i.e.,
(23) $\boldsymbol{X}_{k}^{\mathcal{T}_{k}}=\hat{\boldsymbol{X}}_{k}+\mathbf{T}(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k})),$
where $\mathbf{T}\in\mathbb{R}^{m\times n}$, $\boldsymbol{y}_{k}=\mathcal{H}_{k}(\boldsymbol{x}_{k})$ and $\boldsymbol{\varepsilon}_{k}$ is independent with $\hat{\boldsymbol{X}}_{k}$. Denote
$\bar{\boldsymbol{X}}_{k}$ $=\hat{\boldsymbol{X}}_{k}-\mathbb{E}[\boldsymbol{X}_{k}],$
$\bar{\boldsymbol{Y}}_{k}$ $=\mathcal{H}(\hat{\boldsymbol{X}}_{k})-\boldsymbol{y}_{k}.$
Then the optimization problem Eq. 17 has the global optimal solution
(24) $\mathbf{T}=\mathbb{E}\left[\bar{\boldsymbol{X}}_{k}\bar{\boldsymbol{Y}}_{k}^{\top}\right]\left\{\mathbb{E}\left[\bar{\boldsymbol{Y}}_{k}\bar{\boldsymbol{Y}}_{k}^{\top}\right]+\mathbb{E}[\boldsymbol{\varepsilon}_{k}\boldsymbol{\varepsilon}_{k}^{\top}]\right\}^{-1}.$
Proof 3.7.
By plugging Eq. 23 into Eq. 17, the MMD loss can be rewriten as
$\operatorname{MMD}[p,q;\mathcal{F}]=$ $\mathbb{E}[(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))^{\top}\mathbf{T}^{\top}\mathbf{T}(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))]$
$+2\mathbb{E}[\hat{\boldsymbol{X}}_{k}^{\top}\mathbf{T}(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))]$
$-2\mathbb{E}[\boldsymbol{X}_{k}^{\top}\mathbf{T}(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))]$
@@
$\frac{\partial\mathbf{a}^{T}\boldsymbol{x}\mathbf{b}}{\partial\boldsymbol{x}}=\mathbf{ab}^{T},\qquad\frac{\partial\mathbf{b}^{T}\boldsymbol{x}^{T}\boldsymbol{x}\mathbf{c}}{\partial\boldsymbol{x}}=\boldsymbol{x}\left(\mathbf{b}\mathbf{c}^{T}+\mathbf{cb}^{T}\right),$
@@
$\mathbf{T}\mathbb{E}[(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))^{\top}]=$ $\mathbb{E}[\hat{\boldsymbol{X}}_{k}(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))^{\top}]$
$-\mathbb{E}[{\boldsymbol{X}}_{k}(\boldsymbol{y}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))^{\top}].$
@@
Because $\pi^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$ and $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$ in Eq. 17 are two different distributions, the corresponding random variables $\boldsymbol{X}_{k}^{\mathcal{T}_{k}}$ and $\boldsymbol{X}_{k}$ are set to be independent. Thus
$\mathbf{T}\left\{\mathbb{E}[(\boldsymbol{y}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))(\boldsymbol{y}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))^{\top}]+\mathbb{E}\left[\boldsymbol{\varepsilon}_{k}\boldsymbol{\varepsilon}_{k}^{\top}\right]\right\}=$ $\mathbb{E}[\hat{\boldsymbol{X}}_{k}(\boldsymbol{y}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k}))^{\top}]$
$-\mathbb{E}[{\boldsymbol{X}}_{k}]\mathbb{E}[\boldsymbol{y}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k})]^{\top}.$
Let $\bar{\boldsymbol{X}}_{k}:=\hat{\boldsymbol{X}}_{k}-\mathbb{E}[\boldsymbol{X}_{k}],\bar{\boldsymbol{Y}}_{k}=\mathcal{H}(\hat{\boldsymbol{X}}_{k})-\boldsymbol{y}_{k}$. Then we have
$\mathbf{T}=\mathbb{E}\left[\bar{\boldsymbol{X}}_{k}\bar{\boldsymbol{Y}}_{k}^{\top}\right]\left\{\mathbb{E}\left[\bar{\boldsymbol{Y}}_{k}\bar{\boldsymbol{Y}}_{k}^{\top}\right]+\mathbb{E}[\boldsymbol{\varepsilon}_{k}\boldsymbol{\varepsilon}_{k}^{\top}]\right\}^{-1}.$
Proof is completed.
@@
Since exact states $\boldsymbol{x}_{k}$ and observation $\mathcal{H}_{k}(\boldsymbol{x}_{k})$ are not avaliable, we approximate $\boldsymbol{Y}_{k}$ by $\boldsymbol{y}_{k}^{o}+\boldsymbol{\varepsilon}_{k}$. By Eq. 18, the expectation of posterior $\mathbb{E}\left[\boldsymbol{X}_{k}\right]\approx\sum_{i=1}^{N}w_{i}\hat{\boldsymbol{x}}_{k}^{(i)}$. The ensemble approximation of Eq. 24 are computed by
(25) $\mathbb{E}\left[\bar{\boldsymbol{X}}_{k}\bar{\boldsymbol{Y}}_{k}^{\top}\right]$ $\approx\frac{1}{N-1}\sum_{i=1}^{N}\left(\hat{\boldsymbol{x}}_{k}^{(i)}-\mathbb{E}\left[\boldsymbol{X}_{k}\right]\right)\left(\mathcal{H}(\hat{\boldsymbol{x}}_{k}^{(i)})-\boldsymbol{y}^{o}_{k}\right)^{\top},$
(26) $\mathbb{E}\left[\bar{\boldsymbol{Y}}_{k}\bar{\boldsymbol{Y}}_{k}^{\top}\right]$ $\approx\frac{1}{N-1}\sum_{i=1}^{N}\left(\mathcal{H}(\hat{\boldsymbol{x}}_{k}^{(i)})-\boldsymbol{y}^{o}_{k}\right)\left(\mathcal{H}(\hat{\boldsymbol{x}}_{k}^{(i)})-\boldsymbol{y}^{o}_{k}\right)^{\top},$
(27) $\mathbb{E}[\boldsymbol{\varepsilon}_{k}\boldsymbol{\varepsilon}_{k}^{\top}]$ $\approx\frac{1}{N-1}\sum_{i=1}^{N}\left(\boldsymbol{\varepsilon}^{(i)}_{k}\boldsymbol{\varepsilon}^{(i)}_{k}\right)^{\top},$
where $\{\boldsymbol{\varepsilon}^{(i)}_{k}\}_{i=1}^{N}$ are drawn from observation noise.
Algorithm 2 EnTranF with variance penalty(EnTranFp) on Theorem 3.6
1:Given ensemble $\{\boldsymbol{x}_{k-1}^{(i)}\}_{i=1}^{N}$ at previous time step $k-1$, current observation $\boldsymbol{y}^{o}_{k}$, discrete dynamical model $\mathcal{M}_{k}(\cdot)$, observation operator $\mathcal{H}_{k}(\cdot)$, model uncertainty $\pi(\boldsymbol{\eta}_{k})$ and likelihood function $\pi(\boldsymbol{y}_{k}\mid\boldsymbol{x}_{k})$;
2:for $i=1,\cdots,N$ do
3:      $\hat{\boldsymbol{x}}_{k}^{(i)}=\mathcal{M}_{k}(\boldsymbol{x}_{k-1}^{(i)})+\boldsymbol{\eta}_{k}^{(i)}$;
4:end $\triangleright$ Forecast Step
5:Calculate particle weights $w_{i}=\frac{\pi(\boldsymbol{y}^{o}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(i)})}{\sum_{j=1}^{N}\pi(\boldsymbol{y}^{o}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(j)})}$;
6:Calculate $\mathbf{T}$ in Eq. 23 with ensemble approximation Eq. 25-Eq. 27;
7:for $i=1,\cdots,N$ do
8:      $\boldsymbol{x}_{k}^{(i)}=\hat{\boldsymbol{x}}^{(i)}_{k}+\mathbf{T}(\boldsymbol{y}^{o}_{k}+\boldsymbol{\varepsilon}_{k}-\mathcal{H}(\hat{\boldsymbol{x}}^{(i)}_{k}))$; $\triangleright$ Alternative Analysis Step
9:end for
10:Approximated posterior ensemble $\{{\boldsymbol{x}}_{k}^{(i)}\}_{i=1}^{N}$;

### v2

@@ 3.2.1 Variance penalty for MMD loss
$=\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}^{2}\left\langle\frac{\phi_{\bm{x}}-\phi_{\bm{y}}}{\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}},f\right\rangle_{\mathcal{H}}^{2}\right],$
where $\phi_{\bm{x}}=k(\bm{x},\cdot)$ and $\phi_{\bm{y}}=k(\bm{y},\cdot)$. With Cauchy inequality, it gives
$\left\langle\frac{\phi_{\bm{x}}-\phi_{\bm{y}}}{\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}},f\right\rangle_{\mathcal{H}}\leq\frac{\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}}{\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}}\|f\|_{\mathcal{H}}\leq 1.$
When $f=\frac{\phi_{\bm{x}}-\phi_{\bm{y}}}{\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}}\in\mathcal{F}$, the equality holds. Thus
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\bm{x})-f(\bm{y})\right]^{2}=\mathbb{E}\left[\|\phi_{\bm{x}}-\phi_{\bm{y}}\|_{\mathcal{H}}^{2}\right].$
Then the first term of (3.18) can be rewritten as
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\bm{x})-f(\bm{y})\right]^{2}=\mathbb{E}_{\bm{x}\sim p}[k(\bm{x},\bm{x})]-2\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q}[k(\bm{x},\bm{y})]+\mathbb{E}_{\bm{y}\sim p}[k(\bm{y},\bm{y})].$ (3.19)
The second term is actually the maximum covariance discrepancy (MCD) that proposed in [37]. It can be rewritten as
$\underset{f\in\mathcal{F}}{\sup}\ \left(\mathbb{V}ar_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{V}ar_{\bm{y}\sim q}[f(\bm{y})]\right)=\|\mathbb{C}[p]-\mathbb{C}[q]\|_{\text{HS}},$ (3.20)
where $\|\cdot\|_{\text{HS}}$ denote the Hilbert-Schmidt norm and $\mathbb{C}[p]$ is defined as $\mathbb{C}[p]=\mathbb{E}_{p}[\phi(\bm{x})\otimes\phi(\bm{x})]-\mathbb{E}_{p}[\phi(x)]\otimes\mathbb{E}_{p}[\phi(x)]$ where $\otimes$ is the tensor product in RKHS $\mathcal{H}$. According to Lemma 2 in [37], we have
$\|\mathbb{C}[p]-\mathbb{C}[q]\|_{\text{HS}}^{2}=$ $\mathbb{E}_{\bm{x},\bm{x}^{\prime}}[k^{2}(\bm{x},\bm{x}^{\prime})]-2\mathbb{E}_{\bm{x}}\left[\mathbb{E}^{2}_{\bm{x}^{\prime}}[k(\bm{x},\bm{x}^{\prime})]\right]+\mathbb{E}^{2}_{\bm{x},\bm{x}^{\prime}}[k(\bm{x},\bm{x}^{\prime})]$
$-2\mathbb{E}_{\bm{x},\bm{y}}[k^{2}(\bm{x},\bm{y})]+2\mathbb{E}_{\bm{x}}\left[\mathbb{E}^{2}_{\bm{y}}[k(\bm{x},\bm{y})]\right]$
$+2\mathbb{E}_{\bm{y}}\left[\mathbb{E}^{2}_{\bm{x}}[k(\bm{y},\bm{x})]\right]-2\mathbb{E}^{2}_{\bm{x},\bm{y}}[k(\bm{x},\bm{y})]$
$+\mathbb{E}_{\bm{y},\bm{y}^{\prime}}[k^{2}(\bm{y},\bm{y}^{\prime})]-2\mathbb{E}_{\bm{y}}\left[\mathbb{E}^{2}_{\bm{y}^{\prime}}[k(\bm{y},\bm{y}^{\prime})]\right]+\mathbb{E}^{2}_{\bm{y},\bm{y}^{\prime}}[k(\bm{y},\bm{y}^{\prime})].$
Similar to (3.16), we can estimate (3.19) with ensemble particles, i.e.,
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\bm{x})-f(\bm{y})\right]^{2}=\sum_{i=1}^{m}$ $w_{i}k(\hat{\bm{x}}_{k}^{(i)},\hat{\bm{x}}_{k}^{(i)})-2\sum_{i,j=1}^{m}w_{i}v_{j}k(\hat{\bm{x}}_{k}^{(i)},\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(j)}))$
$+\sum_{i}^{n}v_{i}k(\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(i)}),\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(i)})),$
And the second term of Equation (3.18) can be estimated with
$\|\mathbb{C}[p]-\mathbb{C}[q]\|_{\text{HS}}^{2}=\operatorname{tr}(KWKW),$
where
$K=\left[\begin{matrix}k(\hat{\bm{x}}^{(1)},\hat{\bm{x}}^{(1)})&\cdots&k(\hat{\bm{x}}^{(1)},\hat{\bm{x}}^{(n)})&k(\hat{\bm{x}}^{(1)},\hat{\bm{y}}^{(1)})&\cdots&k(\hat{\bm{x}}^{(1)},\hat{\bm{y}}^{(n)})\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ k(\hat{\bm{x}}^{(n)},\hat{\bm{x}}^{(1)})&\cdots&k(\hat{\bm{x}}^{(n)},\hat{\bm{x}}^{(n)})&k(\hat{\bm{x}}^{(n)},\hat{\bm{y}}^{(1)})&\cdots&k(\hat{\bm{x}}^{(n)},\hat{\bm{y}}^{(n)})\\ k(\hat{\bm{y}}^{(1)},\hat{\bm{x}}^{(1)})&\cdots&k(\hat{\bm{y}}^{(1)},\hat{\bm{x}}^{(n)})&k(\hat{\bm{y}}^{(1)},\hat{\bm{y}}^{(1)})&\cdots&k(\hat{\bm{y}}^{(1)},\hat{\bm{y}}^{(n)})\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ k(\hat{\bm{y}}^{(n)},\hat{\bm{x}}^{(1)})&\cdots&k(\hat{\bm{y}}^{(n)},\hat{\bm{x}}^{(n)})&k(\hat{\bm{y}}_{k}^{(n)},\hat{\bm{y}}^{(1)})&\cdots&k(\hat{\bm{y}}^{(n)},\hat{\bm{y}}^{(n)})\end{matrix}\right],$
and
$W=\left[\begin{matrix}w_{1}-w_{1}^{2}&w_{1}w_{2}&\cdots&w_{1}w_{n}&0&0&\cdots&0\\ w_{2}w_{1}&w_{2}-w_{2}^{2}&\cdots&w_{2}w_{n}&0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ w_{n}w_{1}&w_{n}w_{2}&\cdots&w_{n}-w_{n}^{2}&0&0&\cdots&0\\ 0&0&\cdots&0&v_{1}-v_{1}^{2}&v_{1}v_{2}&\cdots&v_{1}v_{n}\\ 0&0&\cdots&0&v_{2}v_{1}&v_{2}-v_{2}^{2}&\cdots&v_{2}v_{n}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0&v_{n}v_{1}&v_{n}v_{2}&\cdots&v_{n}-v_{n}^{2}\end{matrix}\right].$
Here, $\{\bm{x}^{(i)}\}_{i=1}^{n}$ are weighted samples from $p$ with weights $\{w_{i}\}_{i=1}^{N}$ and $\{\bm{y}^{(i)}\}_{i=1}^{n}$ are weighted samples from $q$ with weights $\{v_{i}\}_{i=1}^{N}$. In filtering, we take $p$ as the reference distribution and $q$ as the approximated distribution given by transport map $\mathcal{T}_{k}$ respectively.
3.2.2 An extension to high dimensional problems
As shown in algorithm 1, the proposed method requires the weighted prior samples to act as the reference posterior. It essentially employs a transport map to replace the resampling step in particle filtering, which enhances the diversity of particles. In slightly higher dimensional problems, the proposed method can reduce the rate at which particle degeneracy occurs. However, the curse of dimensioanlity remains. In high dimensional problems, the particle degeneracy issue in particle filtering is related to the dimensionality of the observation vector. High-dimensional observation vectors can cause even small differences in each component to result in several orders of magnitude difference in particle weights [31].
To address high-dimensional problems, we split the observations into spatially distinct components and apply Bayes’ theorem separately to each component. Sequential processing of batches of observations is a standard technique in data assimilation [39, 40]. We denote the $i$-th component of observation as $\bm{y}_{[i],k}=\mathbf{H}_{[i],k}(\bm{x}_{k})+\epsilon_{[i],k},i=1,\cdots,m$.
Figure 3.3: Schematic of observation split and composition of transport maps generated by each splited observation.
Based on the schematic of observation split shown in panel (a) of Figure 3.3, we obtain
$\pi(\bm{x}_{k},\bm{y}_{k}\mid\bm{y}_{1:k-1})$ $=\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1})\pi(\bm{y}_{k}\mid\bm{x}_{k})$ (3.21)
$=\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1})\prod_{i=1}^{m}\pi(\bm{y}_{[i],k}\mid\bm{x}_{k}).$
We can express (3.21) as a sequence of Bayesian formulas that gradually incorporate observations.
$\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1],k})\propto\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1})\pi(\bm{y}_{[1],k}\mid\bm{x}_{k}),$
$\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:2],k})\propto\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1],k})\pi(\bm{y}_{[2],k}\mid\bm{x}_{k}),$
$\vdots$
$\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:i],k})\propto\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:i-1],k})\pi(\bm{y}_{[i],k}\mid\bm{x}_{k}),$
$\vdots$
$\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:m],k})\propto\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:m-1],k})\pi(\bm{y}_{[m],k}\mid\bm{x}_{k}),$
$\pi(\bm{x}_{k}\mid\bm{y}_{1:k})=\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:m],k}).$
Let $\bm{x}_{k}^{(i)}=\bm{x}_{k}\mid\bm{y}_{1:k-1},\bm{y}_{[1:i],k}$, then $\pi(\bm{x}_{k}^{(i)})$ is the posterior from the likelihood $\pi(\bm{y}_{[i],k}\mid\bm{x}_{k}^{(i-1)})$ and the prior $\pi(\bm{x}_{k}^{(i-1)})$. Based on the proposed method in Section 3.2.1, each step in above sequential process can be alternated by a transport map $\mathcal{T}_{k}^{\bm{y}_{[i],k}}$. Then the transport map from filter prior $\pi(\bm{x}_{k}\mid\bm{y}_{1:k-1})$ to filter posterior $\pi(\bm{x}_{k}\mid\bm{y}_{1:k})$ can be presented as a composition of the series of transport map (panel (b) of Figure 3.3),
$\mathcal{T}_{k}=\mathcal{T}_{k}^{\bm{y}_{[m],k}}\circ\mathcal{T}_{k}^{\bm{y}_{[m-1],k}}\circ\cdots\circ\mathcal{T}_{k}^{\bm{y}_{[1],k}}.$ (3.22)
From the perspective of computational complexity, the proposed observation splitting approach introduces some additional computational costs; However, these are generally acceptable. Firstly, the dimensionality of each individual sub-problem is substantially reduced. Secondly, by converting the problem into multiple low-dimensional observation data assimilation tasks, the required number of particles can be significantly decreased, leading to improved computational efficiency.
3.2.3 Connection with Ensemble Kalman Filter
@@
If we restrict the kernel function to be the linear kernel and the search space of the transport map to be linear, we can obtain an analytical form of the transport map filter. When using a linear kernel function, the RKHS is the space of linear functions, and MMD only matches the means of the distributions. In this case, we can ignore in (3.20) and focus solely on the mean matching in (3.19).
@@
Given the RKHS kernel $k(\bm{x},\bm{x}^{\prime})=\bm{x}^{\top}\bm{x}^{\prime}+1$. Moreover, we assume (3.13) to be a linear transport, i.e.,
$\bm{X}_{k}^{\mathcal{T}_{k}}=\hat{\bm{X}}_{k}+\mathbf{T}(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k})),$ (3.23)
where $\mathbf{T}\in\mathbb{R}^{m\times n}$, $\bm{y}_{k}=\mathbf{H}_{k}(\bm{x}_{k})$ and $\bm{\varepsilon}_{k}$ is independent with $\hat{\bm{X}}_{k}$. Denote
$\bar{\bm{X}}_{k}$ $=\hat{\bm{X}}_{k}-\mathbb{E}[\bm{X}_{k}],$
$\bar{\bm{Y}}_{k}$ $=\mathbf{H}(\hat{\bm{X}}_{k})-\bm{y}_{k}.$
Then the optimization objective (3.19) has the global minimum, if
$\mathbf{T}=\mathbb{E}\left[\bar{\bm{X}}_{k}\bar{\bm{Y}}_{k}^{\top}\right]\left\{\mathbb{E}\left[\bar{\bm{Y}}_{k}\bar{\bm{Y}}_{k}^{\top}\right]+\mathbb{E}[\bm{\varepsilon}_{k}\bm{\varepsilon}_{k}^{\top}]\right\}^{-1}.$ (3.24)
Proof.
By plugging (3.23) into (3.19), the optimization objective can be rewriten as
$\operatorname{MMD}[p,q;\mathcal{F}]=$ $\mathbb{E}[(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))^{\top}\mathbf{T}^{\top}\mathbf{T}(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))]$
$+2\mathbb{E}[\hat{\bm{X}}_{k}^{\top}\mathbf{T}(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))]$
$-2\mathbb{E}[\bm{X}_{k}^{\top}\mathbf{T}(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))]$
@@
$\frac{\partial\mathbf{a}^{T}\bm{x}\mathbf{b}}{\partial\bm{x}}=\mathbf{ab}^{T},\qquad\frac{\partial\mathbf{b}^{T}\bm{x}^{T}\bm{x}\mathbf{c}}{\partial\bm{x}}=\bm{x}\left(\mathbf{b}\mathbf{c}^{T}+\mathbf{cb}^{T}\right),$
@@
$\mathbf{T}\mathbb{E}[(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))^{\top}]=$ $\mathbb{E}[\hat{\bm{X}}_{k}(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))^{\top}]$
$-\mathbb{E}[{\bm{X}}_{k}(\bm{y}_{k}+\bm{\varepsilon}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))^{\top}].$
@@
Because $\pi^{\mathcal{T}_{k}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ and $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ in (3.14) are two different distributions, the corresponding random variables $\bm{X}_{k}^{\mathcal{T}_{k}}$ and $\bm{X}_{k}$ are set to be independent. Thus
$\mathbf{T}\left\{\mathbb{E}[(\bm{y}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))(\bm{y}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))^{\top}]+\mathbb{E}\left[\bm{\varepsilon}_{k}\bm{\varepsilon}_{k}^{\top}\right]\right\}=$ $\mathbb{E}[\hat{\bm{X}}_{k}(\bm{y}_{k}-\mathbf{H}(\hat{\bm{X}}_{k}))^{\top}]$
$-\mathbb{E}[{\bm{X}}_{k}]\mathbb{E}[\bm{y}_{k}-\mathbf{H}(\hat{\bm{X}}_{k})]^{\top}.$
Let $\bar{\bm{X}}_{k}:=\hat{\bm{X}}_{k}-\mathbb{E}[\bm{X}_{k}],\bar{\bm{Y}}_{k}=\mathbf{H}(\hat{\bm{X}}_{k})-\bm{y}_{k}$. Then we have
$\mathbf{T}=\mathbb{E}\left[\bar{\bm{X}}_{k}\bar{\bm{Y}}_{k}^{\top}\right]\left\{\mathbb{E}\left[\bar{\bm{Y}}_{k}\bar{\bm{Y}}_{k}^{\top}\right]+\mathbb{E}[\bm{\varepsilon}_{k}\bm{\varepsilon}_{k}^{\top}]\right\}^{-1}.$
Proof is completed. ∎
@@
Since true states $\bm{x}_{k}$ and $\bm{y}_{k}=\mathbf{H}_{k}(\bm{x}_{k})$ are not avaliable, we replace $\bm{y}_{k}$ by $\bm{y}_{k}^{o}$. By (3.15), the expectation of posterior $\mathbb{E}\left[\bm{X}_{k}\right]\approx\sum_{i=1}^{N}w_{i}\hat{\bm{x}}_{k}^{(i)}$. The ensemble approximation of (3.24) are computed by
$\mathbb{E}\left[\bar{\bm{X}}_{k}\bar{\bm{Y}}_{k}^{\top}\right]$ $\approx\frac{1}{N-1}\sum_{i=1}^{N}\left(\hat{\bm{x}}_{k}^{(i)}-\mathbb{E}\left[\bm{X}_{k}\right]\right)\left(\mathbf{H}(\hat{\bm{x}}_{k}^{(i)})-\bm{y}^{o}_{k}\right)^{\top},$ (3.25)
$\mathbb{E}\left[\bar{\bm{Y}}_{k}\bar{\bm{Y}}_{k}^{\top}\right]$ $\approx\frac{1}{N-1}\sum_{i=1}^{N}\left(\mathbf{H}(\hat{\bm{x}}_{k}^{(i)})-\bm{y}^{o}_{k}\right)\left(\mathbf{H}(\hat{\bm{x}}_{k}^{(i)})-\bm{y}^{o}_{k}\right)^{\top},$ (3.26)
$\mathbb{E}[\bm{\varepsilon}_{k}\bm{\varepsilon}_{k}^{\top}]$ $\approx\frac{1}{N-1}\sum_{i=1}^{N}\left(\bm{\varepsilon}^{(i)}_{k}\bm{\varepsilon}^{(i)}_{k}\right)^{\top}$, (3.27)
where $\{\bm{\varepsilon}^{(i)}_{k}\}_{i=1}^{N}$ are drawn from observation noise.